OpenStudy (anonymous):

A man goes to the bank with $4.00 and asks for change. He is given an equal amount of nickels, dimes and quarters. How many of each is he given?

OpenStudy (unklerhaukus):

\[$4.00 = x\times$0.05+x\times$0.10+x\times$0.25\]

OpenStudy (anonymous):

so how should i answer it

OpenStudy (anonymous):

can you tell me the answer

OpenStudy (unklerhaukus):

factor the x, add the number on the right hand side, then divide to solve for x.

OpenStudy (anonymous):

can you show me how

OpenStudy (anonymous):

can you do it stepp by step

OpenStudy (unklerhaukus):

\[$4.00 = x\times$0.05+x\times$0.10+x\times$0.25\\ $4.00 = x\times\Big($0.05+$0.10+$0.25\Big)\]

OpenStudy (anonymous):

so what is next can you please tell me the final answer by showing me step by step

OpenStudy (unklerhaukus):

Now, add the numbers on the right hand side,

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i mean after that]

OpenStudy (unklerhaukus):

\[$0.05+$0.10+$0.25\]

OpenStudy (anonymous):

oh i got it thanks

OpenStudy (anonymous):

0.05+$0.10+$0.25 whats the answer

OpenStudy (unklerhaukus):

you got it?

OpenStudy (anonymous):

yes thanks

OpenStudy (unklerhaukus):

how many coins?

OpenStudy (anonymous):

i just have a question how do you divde 4 by 0.4

OpenStudy (anonymous):

i just have a question how do you divde 4 by 0.4

OpenStudy (anonymous):

please tell me

OpenStudy (unklerhaukus):

4/0.4 = 40/4 =

OpenStudy (anonymous):

so the answer is 10 dimes,10 nickels and 10 quaters

OpenStudy (anonymous):

right

OpenStudy (unklerhaukus):

yep

OpenStudy (anonymous):

Thank you so much